Part C: WACC and Capital Budgeting Calculate the firm’s WACC (using 2018 numbers). (You will need to collect information on the long-term debt and common stock equity from the Balance Sheet. The firm has no preferred stock). Use the WACC to calculate NPV and evaluate IRR for proposed capital budgeting projects. Assume the projects are mutually exclusive and the firm has the money available to fund the project.

Part D: Analysis You will must offer suggests to a senior financial manager and CFO on the proposed projects. Be sure to include a discussion of external funding and where it should come from if necessary and which project the firm should undertake.

Project A Project B Project C Year -100,000 250,000 -500,000 Year 1 50,000 100,000 400,000 Year 2 30,000 100,000 50,000 Year12/31/18 12/31/17 12/31/16 12/31/15 Current Assets Cash And Cash Equivalents Short Term Investments Net Receivables InventoryΕ 12/31/15 12/31/18 12/31/17 12/31/16 2 Revenue 3 Total Revenue 4 Cost of Revenue 5 Gross Profit 51,728,000 57,902,000 56,519​​​​​​​
Answer #1

Answer:

WACC = wD*rD(1-t) + wP*rP + wE*rE
where,
w = the respective weight of debt, preferred stock/ equity and the equity in total capital structure
t = tax rate
D = cost of debt
P = cost of preferred stock/ equity
E = cost of equity

WACC = 0.606(0.0431)(1-.40) + 0.394 (0.2165) = 10.10 %  

= 10% (approx)*

Working Notes:

Income tax rate calculation:
Income tax expense 2,696,000
Income before tax

6,674,000

rate = Income tax expense/income before tax

= (2,696,000/6,674,000)*100

= 40 % (approx)*

* Note - Approx values have been considered for easier calculation.

Calculation of Rate of Debt:

= Interest Expense/ Long term debt = 1,220,000/ 28,293,000 = 4.31%

Calculation of Rate of Equity:

= Net Income / Equity = 3978,000 / 18375,000 = 21.65 %

Calculation of Weights:

Long term debt = 28,293,000 : Weight of Debt (wD) = 28293,000/46,668,000 = 0.606

Share holder's Equity = 18,375,000 : Weight of Equity (wE) = 18,375,000/46,668,000 = 0.394

Total Capital = Debt + Equity = $ 46,668,000

Calculation of NPV:

Discount Rate 10%
Year 0 1 2 3 4
Project A                       (100,000)                          50,000                          30,000                      30,000                      20,000
Project B                       (250,000)                       100,000                       100,000                      50,000                      50,000
Project C                       (500,000)                       400,000                          50,000                      50,000                      50,000
NPV for Project A $5,861.50
NPV for Project B ($4,299.88)
NPV for Project C ($21,204.46)